### v1

# Graphs and Their Vietoris-Rips Complexes Have the Same Pseudotopological Weak Homotopy Type

Abstract.
In this document, we propose a bridge between the graphs and the geometric realizations of their Vietoris Rips complexes, i.e. Graphs, with their canonical Čech closure structure, have the same homotopy type that the realization of their Vietoris Rips complex.
This work was supported by the CONACYT postgraduate studies scholarship number 839062, and the results in this article form a part of the author’s PhD thesis project supervised by Antonio Rieser.
1. Introduction
The homotopical study of (reflexive) graphs has recently seen increased activity and interest from a number of different perspectives. For example, in [Babson_etal_2006] the authors develop $A$-theory, a cubical-type homotopy is defined in graphs and uses the inductive product and the discrete interval to generate their homotopies; in [Dochtermann_2009, Chih_Scull_2021] $\times$-homotopy is introduced for studying not-necessarily reflexive graphs; there are also homotopy theories for digraphs [Grigoryan_etal_2014] and the non-topological homotopy for metric spaces considered in [Plaut_Wilkins_2013]; restricting to reflexive graphs on cubical lattices whose edges are detemined by their position in the lattice, we obtain digital homotopy [Lupton_etal_2022a, Lupton_Scoville_2022], finally, Antonio Rieser proposed that Čech closure spaces [Rieser_2021] and pseudotopological spaces [Rieser_arXiv_2022] could be used as categories in which to simultaneously develop the homotopy theory of topological spaces, graphs, digraphs, and point clouds (the latter for topological data analysis [Carlsson_2009, Carlsson_et_al_2012]), extending earlier work of Demaria [Demaria_1987] on graphs.
In this article, we continue the development of homotopy theory in pseudotopological spaces begun in [Rieser_2021, Rieser_arXiv_2022], and we prove that every graph $X$ has the same weak homotopy type as its Vietoris-Rips complex $|VR(X)|$. To achieve this, we define an intermediate pretopological space $R(X)$ (2.6), which is a simplicial complex with a modified closure structure such that (1) there is a continuous inclusion of the vertices of $X$ into $R(X)$, (2) $R(X)$ is strong homotopy equivalent to the graph, and (3) $R(X)$ is weak homotopy equivalent to the geometric realization $|VR(X)|$ of the Vietoris-Rips complex of $X$. The construction of the weak homotopy equivalence further requires that we construct three transformations of functions $f:Y\to R(X)$, which are the key to the proof:
  • the discrete modification (2.10), which sends a map $f$ with codomain $R(X)$ to one whose image lies only in $X\subset R(X)$;
  • the flood (2.15), which expands the preimage of a vertex $v\in X\subset R(X)$ in order to create a neighorhood $U$ of $f^{-1}(v)\subset Y$ small enough so that the image of every element in $U$ is contained in the closure of the image of any of its points;
  • the convex transformation (2.20), which converts maps $f:Y\to R(X)$ whose image is in $X\subset R(X)$ into a simplicial map from a triangulation of $Y$ to $R(X)$.
We study the properties of each of these transformations in turn, which we then use in combination for the proof of our main result. The critical property of each is that these transformations, when applied to a continuous maps from a triangulizable compact metric space, produce a continuous map.
Applying the flood transformation after the discrete modification uses the same fundamental idea as that studied in digital homotopy to find the second homotopy group for the digital sphere [Lupton2023DigitalHom]. In contrast to that paper, whose domain is the lattice $\mathbb{Z}^{2}$ and the distances are discrete, in this paper, the domain will be a triangulizable compact metric space, in particular $S^{n}$ and $S^{n}\times I$.
The results in this paper, when combined with well-known results about the Vietoris-Rips complex [Latschev_2001], we observe that for a “big enough” cloud of points sampled from a manifold, the pseudotopological homotopy groups (defined in [Rieser_arXiv_2022]) from certain closure spaces built on those points are isomorphic to the classical homotopy groups of the manifold being sampled. In addition, the Vietoris-Rips complex is difficult to understand and only a few complete calculations are known for metric spaces at any scale, for example [Adams2023lower, Adams2022VietorisRipsHypercube, Adams2020VietorisRipsMetricGluings] and is used to bound Gromov-Hausdorff distance [Adams2022gromovhausdorff, Memoli2023GHUltraMetricSpaces, Memoli2023GHSpheres]; thus, with this paper and growing wave of studies in these categories mentioned above, we may hope for new ways to approach the Vietoris-Rips complex as well.
The isomorphism between the pseudotopological homotopy groups of finite graphs and those of its clique complex has been also established independently by Milicevic and Scoville in [WebJMMNicks]. Their proof is very different from this one, particularly as it does not pass through the space $R(X)$, and it only applies to finite graphs, whereas the technique we study here is also valid for arbitrary, even uncountably infinite, reflexive graphs.
2. Weak Homotopy Equivalence
Closure spaces have two definitions which are equivalent categories: one defined through the closure operation, which is a map from the power set to the power set, and another defined through convergence of filters, which is a function from the space to the collection of all of the filters. For this paper, it is enough to know that we can use both definition for the same object. By completeness, we briefly introduce these both categories. (To go deeper with why these are equivalent, we invite the reader to check [Preuss_2002], proposition 2.3.1.8)
Definition 2.1 ([Preuss_2002], 2.3.1.6).
A closure space is a pair $(X,c)$, where $X$ is a set and $c:2^{X}\rightarrow 2^{X}$ a function, called a closure operation on $X$, satisfying the following conditions:
  1. (1)
    $c(\varnothing)=\varnothing$.
  2. (2)
    $A\subset c(A)$ for each $A\in 2^{X}$.
  3. (3)
    $c(A\cup B)=c(A)\cup c(B)$ for each $A,B\in 2^{X}$.
A map $f:X\rightarrow Y$ is continuous from $(X,c)$ to $(Y,c^{\prime})$, closure spaces, denoted as $f:(X,c)\rightarrow(Y,c^{\prime})$, provided that $f(c(A))\subset c^{\prime}(f(A))$ for every $A\in 2^{X}$.
For any collection $\mathcal{A}\subset 2^{X}$ such any finite intersection is non-empty, we define the filter generated by $\mathcal{A}$ on $X$, $[\mathcal{A}]_{X}$, as the collection of $B\in 2^{X}$ such that either $B=X$ or for $B\supset A_{1}\cap\ldots\cap A_{n}$ for some $A_{i}\in\mathcal{A}$. We denote the filter generated by $\mathcal{A}=\{A\}$ as $[A]_{X}$ and the filter generated by $A=\{x\}$ as $[x]_{X}$; when the context is clear, we omit the set which the filter is on. For a function $f:X\rightarrow Y$ and $\mathcal{F}$ a filter on $X$, $f(\mathcal{F})\coloneqq[f(A)\mid A\in\mathcal{F}]_{Y}$ is said to be the image of $\mathcal{F}$ under the map $f$.
Remark.
Let $\mathcal{A},\mathcal{B}\subset 2^{X}$ closed under finite intersections. If for every $B\in\mathcal{B}$ there exists $A\in\mathcal{A}$ such that $A\subset B$, then $[\mathcal{B}]_{X}\subset[\mathcal{A}]_{X}$. In particular, if $A\subset B$, then $[B]_{X}\subset[A]_{X}$.
Definition 2.2 ([Cech_1966], 14 B.1).
Let $(X,c)$ a closure space. We say that $U\subset X$ is neighborhood of a point $x\in X$ if $x\in X\backslash c(X\backslash U)$. The neighborhood system of a point $x$, denoted $\mathcal{U}(x)$, is the collection of neighborhoods of $x$.
Proposition 2.3 ([Preuss_2002], 2.3.1.6).
Let $(X,c)$ and $(Y,c^{\prime})$ closure spaces. Then $f:X\rightarrow Y$ is continuous if and only if $f(\mathcal{U}(x))\supset\mathcal{U}(f(x))$ for every $x\in X$.
Example.
Let $S^{1}$ be the circle (of diameter one) and $C_{4}$ be 4-cyclic graph (or the digital circle), i.e. the set $\{0,1,2,3\}$ such that $cl(y)=\{y-1,y,y+1\}$ module 4. We have that the map $f:S^{1}\rightarrow C_{4}$ such that
$f(x)=\left\{\begin{array}[]{ll}0&x\in[0,1/4)\\ 1&x\in[1/4,2/4)\\ 2&x\in[2/4,3/4)\\ 3&x\in[3/4,1)\end{array}\right.$
is continuous.
Definition 2.4.
Let $f,g:X\rightarrow Y$ be continuous maps between closure spaces. $f$ and $g$ are said to be homotopic, denoted by $f\simeq g$, if there exists a continuous map $H:X\times I\rightarrow Y$ such that $H(x,0)=f(x)$ and $H(x,1)=g(x)$ for every $x\in X$. $X$ and $Y$ are called strong homotopy equivalent if there exist continuous maps $f:X\rightarrow Y$ and $g:Y\rightarrow X$ such that $gf\simeq 1_{X}$ and $fg\simeq 1_{Y}$.
We define the homotopy classes for a closure space $X$ for every $n\geq 0$, denoted by $\pi_{n}(X,*)$, as the set of equivalence classes of continuous maps $(S^{n},*)\rightarrow(X,*)$. Using the standard $\star$ operation in those maps and standard arguments, we can prove that $\pi_{n}(X,*)$ is a group for all $n\geq 1$ and an abelian group for every $n>1$. Two closure spaces $X$ and $Y$ are weak homotopy equivalent, and $f$ (respectively $g$) is a weak homotopy equivalence if there exists a continuous map $f:X\rightarrow Y$ (or $g:Y\rightarrow X$) such that $\pi_{n}(f):\pi_{n}(X,*)\rightarrow\pi_{n}(Y,*)$ (or $\pi_{n}(g):\pi_{n}(Y,*)\rightarrow\pi_{n}(X,*)$) is an isomorphism for every $n\geq 0$. We say that $X$ and $Y$ have the same weak homotopy type if there exists a zig-zag of weak homotopy equivalences from $X$ to $Y$
$X\rightarrow X_{1}\leftarrow X_{2}\rightarrow\ldots\leftarrow X_{n-1}\rightarrow X_{n}\leftarrow Y.$
Remark.
A Čech closure space is a particular case of a pseudotopological space. The pseudotopological weak homotopy studied in [Rieser_arXiv_2022] agrees with the one defined for closure spaces in 2.4.
Every undirected (also valid for directed) simple graph has a natural Čech closure structure and induces a simplicial complex. We recall these structures in the following definition.
Definition 2.5.
Let $(X,E)$ be a simple, reflexive graph (i.e. the graph does not have multiple edges and every loop $(x,x)$ is an edge in the graph). We say that $(X,E)$ is locally finite if for every $x\in X$ there exists a finite number of vertices $(x,y)\in E$. We define the canonical closure structure for a graph $(X,E)$, denoted by $c_{X}$, such that
$c_{X}(x)\coloneqq$$\{y\in X\mid(x,y)\in E\}\text{ for every }x\in X$
$c_{X}(A)\coloneqq$ $\bigcup\limits_{x\in A}c_{X}(A)\text{ for every }A\subset X$
Observe that this closure operation induces the neighborhood system at $x$ in $X$ as $\mathcal{U}_{X}(x)=[c_{X}(x)]_{X}$.
When the context is clear, we omit to write the set of edges, i.e. we write just $X$ instead of $(X,E)$ when we refer to a graph. We can (and we implicitly do) set an arbitrary order $<$ over the vertices of $X$; this is helpful in a couple of construction during the document.
The following definition introduce two realizations of a graph. The first realization is the classical simplicial complex induced by a graph (the clique complex); on the other hand, the second realization is a hybrid between the graph and the clique complex. This second realization is actually finer than the first one and allows a strong homotopy equivalence between itself and the graph.
Definition 2.6.
Let $X$ be a graph.
  1. (1)
    $|VR(X)|$ is the geometric realization of the simplicial complex with $n$-simplices $\{v_{0},\ldots,v_{n}\}$ such that $v_{i}\in cl(v_{j})$ for every $i,j$. This space is a topological, and then it has a natural topological closure operation which we denote by $c_{|VR(X)|}$.
  2. (2)
    $R(X)$ is defined over the same set of $|VR(X)|$, but the neighborhood at $v\in X$ in $R(X)$ is defined as
    $\mathcal{U}_{v}\coloneqq\left[\bigcup\limits_{v\in\sigma}|\sigma|\right]_{R(X)}$
    which is a neighborhood system at $v$. The neighborhood system of every point in $|VR(X)|-X$ is the equal to the neighborhood filter in $|VR(X)|$. We denote the closure defined by the neighborhood system as $c_{R(X)}$.
It is straightforward to see that $(|VR(X)|,c_{|VR(X)|})$ is a finer pretopological space than $(|VR(X)|,c_{R(X)})$, they are over the same set and by construction there are fewer sets in the neighborhood at $v$ in $R(X)$ than in $|VR(X)|$ (i.e. the identity is continuous as PsTop spaces, and then as closure spaces). In the same sense, we construct $R(X)$ satisfying that $X$ is a (pretopological) subspaces of it.
Definition 2.7 (Relative Interior of a Simplex).
Let $\Delta=\{v_{0},\ldots,v_{n}\}$ be a $n$-simplex. We define the boundary of $|\Delta|$ as the set
$\partial|\Delta|\coloneqq\bigcup_{i\in[n]}\{v_{0},\ldots,\hat{v}_{i}\ldots,v_{n}\},$
with $\{v_{0},\ldots,\hat{v}_{i}\ldots,v_{n}\}$ being the set $\Delta$ excluding $v_{i}$, and we define the relative interior of $\Delta$ as $int(\Delta)=|\Delta|\backslash\partial|\Delta|$.
Definition 2.8 (The Barycentric Cover of a Simplex).
Let $\Delta^{n}\coloneqq\{v_{0},\ldots,v_{n}\}$ an $n$-simplex. Define
$b_{\{v_{i_{0}},\dots,v_{i_{k}}\}}\text{ the barycenter of the $k$-simplex }\{v_{i_{0}},\dots,v_{i_{k}}\}$
for every $i\in[n]$, let’s define
$A_{i}\coloneqq\bigcup_{\alpha_{j}\in[n]}|\{v_{i},b_{\{v_{i},v_{\alpha_{1}}\}},\ldots,b_{\{v_{i},v_{\alpha_{1}},\dots,v_{\alpha_{n}}\}}\}|$
The collection $\{A_{0},\ldots,A_{n}\}$ is said to be the barycentric cover of $|\Delta^{n}|$
Proposition 2.9.
Let $\Delta^{n}\coloneqq\{v_{0},\ldots,v_{n}\}$ be an $n$-simplex and $\{A_{0},\ldots,A_{n}\}$ the barycentric cover of $\Delta^{n}$. Then $|\Delta^{n}|=\cup_{i=0}^{n}|A_{i}|$.
Proof.
Let $x\in|\Delta^{n}|$, then for every $i\in[n]$ there exists $t_{i}\geq 0$ such that $\sum_{i=0}^{n}=1$ and
$x=\sum_{i=0}^{n}t_{i}v_{i}$
Since we have a finite amount of elements, there exists and index in $[n]$, say $\beta_{0}$ such that $t_{\beta_{0}}=\min_{i\in[n]}\{t_{i}\}$, thus
$x=\sum_{i\in[n]\backslash\{\beta_{0}\}}(t_{i}-t_{\beta_{0}})x_{i}+\frac{t_{\beta_{0}}}{n+1}\left(\frac{1}{n+1}\sum_{i=0}^{n}x_{i}\right)$
With the same argument we can find the index $\beta_{1}$ such that $t_{\beta_{1}}=\min_{i\in[n]\backslash\beta_{0}}$ and
$x=$ $\sum_{i\in[n]\backslash\{\beta_{0},\beta_{1}\}}(t_{i}-t_{\beta_{1}})x_{i}+(t_{\beta_{1}}-t_{\beta_{0}})n\left(\frac{1}{n}\sum_{i\in[n]\backslash\{\beta_{0}\}}x_{i}\right)$
$+t_{\beta_{0}}(n+1)\left(\frac{1}{n+1}\sum_{i=0}^{n}x_{i}\right)$
Following this procedure until we cover all of the vertexes, we obtain that
$x=\sum_{k=0}^{n}\left((t_{\beta_{n-k}}-t_{\beta_{n-k-1}})(k+1)\left(\frac{1}{k+1}\sum_{i\in[n]\backslash\{\beta_{j}\mid j\in[k]\}}x_{i}\right)\right)$
defining $t_{\beta_{-1}}=0$. Observe that, by definition of barycenter,
$b_{\{v_{\beta_{j}}\mid j\in[k]\}}=\frac{1}{k+1}\sum_{i\in[n]\backslash\{\beta_{j}\mid j\in[k]\}}x_{i}$
Concluding that $x\in A_{\beta_{n}}$. ∎
Take a $n$-simplex $\Delta^{n}\coloneqq\{v_{0},\ldots,v_{n}\}$ and a $m$-simplex $\Delta^{m}\coloneqq\{w_{0},\ldots,w_{m}\}$ which share a $k$-simplex, say $\Delta^{k}\coloneqq\{v_{0},\dots,v_{k}\}$ with $v_{i}=w_{i}$ for every $i\in[k]$; suppose that $\{A_{0},\ldots,A_{n}\}$, $\{B_{0},\ldots,B_{m}\}$ and $\{C_{0},\ldots,C_{k}\}$ the barycentric cover of $\Delta^{n}$, $\Delta^{m}$ and $\Delta^{k}$, respectively. Observe that $C_{i}=A_{i}\cap\Delta^{n}=B_{i}\cap\Delta^{m}$ for every $i\in[k]$ given that the barycentric cover of $C_{i}$ only depends on the barycenters in $\Delta^{k}$.
With this in mind, we can extend the definition below (2.10) for all $R(X)$ if we can define it in the simplexes.
Definition 2.10 (Discrete Modification).
Let $X$ be a graph, we define the discrete modification as $\Theta_{X}:R(X)\rightarrow X\subset R(X)$ as follows: Let $\Delta^{n}\coloneqq\{v_{0},\ldots,v_{n}\}$ be an $n$-simplex and define
$b_{\{v_{i_{0}},\dots,v_{i_{k}}\}}\text{ the barycenter of the $k$-simplex }\{v_{i_{0}},\dots,v_{i_{k}}\}$
for every $i\in[n]$, recall that
$A_{i}\coloneqq\bigcup_{\alpha_{j}\in[n]}|\{v_{i},b_{\{v_{i},v_{\alpha_{1}}\}},\ldots,b_{\{v_{i},v_{\alpha_{1}},\dots,v_{\alpha_{n}}\}}\}|$
Observe that if we have some repeat elements, then they are just simplexes contained in the one which every element is different. Then we define:
$\Theta_{X}(x)\coloneqq$$\left\{\begin{array}[]{ll}v_{0}&x\in A_{0}\\ v_{1}&x\in A_{1}-A_{0}\\ \vdots\\ v_{n}&x\in A_{n}-\cup_{i=0}^{n-1}A_{i}\end{array}\right.$
Figure 1. Illustration of the discrete modification for the geometric realization of a 2-simplex
Lemma 2.11.
Let $X$ be a graph, the discrete modification $\Theta_{X}:R(X)\rightarrow X\subset R(X)$ is continuous.
Proof.
Let $x\in R(X)$. In the first case, let $x\in X$, by definition we have that $\mathcal{U}(x)=[\cup_{x\in\sigma}|\sigma|]_{R(X)}$. By the definition of the discrete modification, $\Theta_{X}|\sigma|=\{v\mid v\in\sigma\}\subset|\sigma|$. Then $\Theta_{X}(\cup_{x\in\sigma}|\sigma|)=\cup_{x\in\sigma}\{v\mid v\in\sigma\}\subset\cup_{x\in\sigma}|\sigma|$. Thus $\Theta_{X}[\cup_{x\in\sigma}|\sigma|]_{R(X)}\rightarrow x=\Theta_{X}(x)$.
Let $x\notin X$, by the definition of $VR(X)$, there exists a unique simplex $\sigma$ containing $x$ in its relative interior, then $\Theta_{X}(x)\in\{v\mid v\in\sigma\}$. A filter $\mathcal{F}\rightarrow x$ in $VR(X)$ (and then in $R(X)$ because $x\notin X$) if it converges in the restriction of every simplex which contains $x$, then $\cup_{\{v\mid v\in\sigma\}\subset\sigma^{\prime}}|\sigma^{\prime}|$. Then $\Theta_{X}(\mathcal{U}(x))\supset\Theta_{X}[\cup_{\{v\mid v\in\sigma\}\subset\sigma^{\prime}}|\sigma^{\prime}|]_{R(X)}\supset[\cup_{\Theta_{X}(x)\in\sigma}|\sigma|]_{R(X)}$. Thus $\Theta_{X}(\mathcal{U}(x))\rightarrow\Theta_{X}(x)$. ∎
Remark.
By construction, we have that $Im(\Theta_{X})\subset X$, $\Theta_{X}\Theta_{X}=\Theta_{X}$ and $\Theta_{X}(v)=v$ for every $v\in X\subset R(X)$.
When $(X,d)$ is a metric space, we write $B_{r}(x)\coloneqq\{y\in X\mid d(x,y)<r\}$.
Proposition 2.12.
Let $X$ be a graph. $X$ and $R(X)$ are strong homotopy equivalent.
Proof.
Let $g:X\rightarrow R(X)$ be the map which sends the points in $X$ to the corresponding vertices of the simplices in $R(X)$.
It is straightforward to observe that $\Theta_{X}g=1_{X}$, so we only need to express a homotopy from $g\Theta_{X}$ to $1_{R(X)}$.
First we restrict for a simplex $\Delta\in VR(X)$. Let $H\mid_{\Delta}:\Delta\times I\rightarrow\Delta\subset R(X)$ be defined by
$H\mid_{\Delta}(x,t)\coloneqq\left\{\begin{array}[]{ll}1_{\Delta}(x)&\text{ if }t\in[0,1)\\ \Theta_{X}(x)&\text{ if }t=1.\end{array}\right.$
We claim that $H\mid_{\Delta}$ is continuous: For every point $(x,1)$, its image contained in $\Delta$, getting that $H$ is continuous in that point; for every point $(x,t)$ with $t<1$, there exist $r>0$ such that $B_{r}(x,t)\cap(\Delta\times\{1\})=\varnothing$, concluding that $H(B_{r}(x,t))=B_{r}(x)$ and $H$ is continuous in that point.
Denote the neighborhood system of $t$ in $[0,1]$ by $\mathcal{I}(t)$. For the whole domain $R(X)$, we observe that for all $t<1$ there exists $\varepsilon>0$ such that for every $U\in\mathcal{U}$ satisfies that $U\times(t-\varepsilon,t+\varepsilon)\cap U\times\{1\}=\varnothing$, then $H(\mathcal{U}(x)\times(t-\varepsilon,t+\varepsilon))=\mathcal{U}(x)$, and then $H(\mathcal{U}(x)\times\mathcal{I}(t))\rightarrow x$.
On the other hand, $H(\mathcal{U}\times\mathcal{I}(1))=[\{H(A\times[1-1/n,1)\mid A\in\mathcal{U}(x),n\in\mathbb{N}\}]_{R(X)}=\{\Theta_{X}(A)\cup U(A)\mid A\in\mathcal{U}(x)\}]_{R(X)}$ by definition of $H$. As we see in 2.11, there exists an element $B\in\mathcal{U}(x)$ such that $\Theta_{X}(B)\subset\cup_{\Theta_{X}(x)\in\sigma}|\sigma|$ and $B\subset\cup_{\Theta_{X}(x)\in\sigma}|\sigma|$ for every $B\subset A$. Thus, $H(\mathcal{U}(x)\times\mathcal{I}(1))\subset[\cup_{\Theta_{X}(x)\in\sigma}|\sigma|]$ concluding that $H(\mathcal{U}(x)\times\mathcal{I}(1))\rightarrow\Theta_{X}(x)$. ∎
Lemma 2.13.
Let $X$ be a graph, $Y$ be a compact metric space with the topology induced by the metric and suppose that $f:Y\rightarrow R(X)$ is continuous. Then $f_{d}\coloneqq\Theta_{X}f:Y\rightarrow R(X)$ is a continuous map such that
  1. (1)
    $Im(f_{d})\subset X\subset R(X)$,
  2. (2)
    If $f(y)\in X\subset R(X)$, then $f_{d}(y)=f(y)$,
  3. (3)
    $f_{d}\simeq f$.
In addition, if $X$ is locally finite, then $f$ satisfies that
  1. (4)
    #$\#(Im(f_{d}))<\infty$
Proof.
First suppose that $X$ is any graph. Let $f:Y\rightarrow R(X)$ a morphism in PreTop, and define $f_{d}\coloneqq\Theta_{X}f$. Then by definition $Im(f_{d})\subset X$ and $f_{d}(y)=f(y)$ if $f(y)\in X$.
To observe that $f_{d}\simeq f$, 2.12 establishes that $\Theta_{X}$ is homotopic to the identity $1_{R(X)}:R(X)\rightarrow R(X)$. Thus, $f_{d}=\Theta_{X}f\simeq 1_{R(X)}f=f$.
In the last part of the proof, suppose that $X$ is locally finite. Since $f_{d}$ is continuous, being the composition of continuous maps, and $cl(f_{d}(y))$ is the minimal neighborhood, then for every $y\in Y$ there exists $r_{y}>0$ such that $f_{d}(B_{r_{y}}(y))\subset cl(f_{d}(y))$ since $cl(f_{d}(y))$ is a neighborhood of $f_{d}(y)$ in $R(X)$. Given that $X$ is locally finite, then #$\#f_{d}(B_{r_{y}}(y))\leq\#(cl(f_{d}(y))\cap X)<\infty$. Since $Y$ is compact, there exists a finite number of points of $Y$, say $\{y_{1},\ldots,y_{k}\}$, such that $Y=\cup_{i=1}^{k}B_{r_{y_{i}}}(y_{i})$. Thus #$\#f_{d}(Y)\leq\sum_{i=1}^{k}\#f_{d}(B_{r_{y_{i}}}(y_{i}))$, i.e. #$\#f_{d}(Y)$ is finite. ∎
The following lemma removes the sufficient condition in 2.13 that $X$ be a locally finite graph.
Lemma 2.14.
Let $X$ be a graph, $Y$ be a compact metric space with the topology induced by the metric, $A\subset Y$ a compact subset, $v\in X$, and suppose that $f:(Y,A)\rightarrow(R(X),v)$ be continuous. Then there exists $f^{\prime}_{d}:(Y,A)\rightarrow(X,v)$ a continuous map such that
  1. (1)
    $Im(f^{\prime}_{d})\subset X\subset R(X)$.
  2. (2)
    #$\#(Im(f^{\prime}_{d}))<\infty$.
  3. (3)
    $f^{\prime}_{d}(A)=v$.
  4. (4)
    $f^{\prime}_{d}\simeq f$.
Proof.
Define $f_{d}\coloneqq\Theta_{X}f:Y\rightarrow R(X)$. By 2.13, we obtain that
  • $Im(f_{d})\subset X\subset R(X)$.
  • $f_{d}(A)=v$, because the one hypothesis is that $f(A)=v$.
  • $f_{d}\simeq f$.
The following construction is very similar to the subsequent proof of 2.17 but with some nuances: Since for every $x\in X$, $cl(x)$ is the minimal neighborhood of $x$, so for all $y\in Y$ there exists $r_{y}>0$ such that $f_{d}(B_{r_{y}}(y))\subset cl(f_{d}(y))$. In addition, since $A$ is a compact subset of a topological Hausdorff space by hypothesis, $A$ is closed, so we can choose $r_{y}$ sufficiently small so that $B_{r_{y}}(y)\cap A=\varnothing$ for every $y\in Y\backslash A$. Then $\{B_{r_{y}/2}(y)\mid y\in Y\}$ is a open cover of the compact space $Y$, and there exists a finite number of points $\{y_{1},\ldots,y_{n}\}$ such that
$Y=\bigcup\limits_{i=1}^{n}B_{r_{y_{i}}/2}(y_{i})$
We need two additional observations before we can construct a candidate for $f^{\prime}_{d}$. First, observe that if $y_{i}\notin A$, then $B_{y_{i}/2}(y_{i})\subset B_{y_{i}}(y_{i})\subset Y\backslash A$. Second, consider $y_{i}$ and $y_{j}$ such that $B_{r_{y_{i}}/2}(y_{i})\cap B_{r_{y_{j}}/2}(y_{j})\neq\varnothing$, then
$d(y_{i},y_{j})<(r_{y_{i}}+r_{y_{j}})/2\leq\max(r_{y_{i}},r_{y_{j}})$
and thus $y_{i}\in B_{r_{y_{j}}}(y_{j})$ or $y_{j}\in B_{r_{y_{i}}}(y_{i})$; both imply that $f_{d}(y_{i})\in cl(f_{d}(y_{j}))$, recalling that the closure structure on $X\subset R(X)$ is symmetric (i.e $y_{i},y_{j}\in X\subset R(X)$ and $y_{i}\in cl(y_{j})\iff y_{j}\in cl(y_{i}))$), since the closure structure on $X$ is induced by a graph.
For every $y_{i}\in\{y_{1},\dots,y_{n}\}$, define $f_{i}:Y\to X\subset R(X)$ by
$f_{i}(y)\coloneqq\begin{cases}f_{d}(y_{i})&y\in B_{r_{y_{i}}/2}(y_{i}),\\ f_{i-1}(y)&\text{otherwise},\end{cases}$
where $f_{0}=f_{d}$. We first note that, by definition, since $Im(f_{d})\subset X\subset R(X)$, it follows that $Im(f_{i})\subset X\subset R(X)$ for every $i\in\{1,\dots,n\}$. We now claim that, for every $i\in\{1,\dots,n\}$, each $f_{i}$ is continuous, $f_{i}\simeq f_{i-1}$, and $f_{i}(A)=v$.
By the observations above, $B_{r_{y_{i}}/2}(y_{i})\subset Y\backslash A$ for any $y_{i}\in\{y_{1},\dots,y_{n}\}$ with $y_{i}\in Y\backslash A$, and therefore for every such $y_{i}\in Y\backslash A$, $f_{i}(y)=f_{i-1}(y)$ for any point $y\in A$. Also note that for any $y_{i}\in A$, $f_{i}(y_{i})=v$. Since, by construction, $A$ is covered by the $B_{r_{y_{i}}}(y_{i})$ with $y_{i}\in A$, we conclude that $f_{i}(A)=v$.
We now prove that $f_{i}$ is continuous, by proceeding by induction over $i$. Observe first that $f_{0}$ is continuous by 2.13, we obtain the base case. Suppose that $f_{i-1}$ is continuous and consider the following three cases:
  1. Case 1:
    $y\in B_{r_{y_{i}}/2}(y_{i})$. Then there exists $r>0$ such that $B_{r}(y)\subset B_{r_{y_{i}}/2}(y_{i})$, and then $f_{i}(y)=f_{i}(y_{i})=f_{d}(y_{i})$, and, in particular, $f_{i}(B_{r}(y))=f_{i}(B_{r_{y_{i}}/2}(y_{i}))=\{f_{i}(y)\}\subset cl(f_{i}(y_{i}))=cl(f_{i}(y))$.
  2. Case 2:
    $y\in\partial B_{r_{y_{i}}/2}(y_{i})$. By the induction hypothesis, $f_{i-1}$ is continuous, and therefore there exists $r>0$ such that and $B_{r}(y)\subset B_{r_{y_{i}}}(y_{i})$ and $f_{i-1}(B_{r}(y))\subset cl(f_{i-1}(y))$.
    Now define $\Lambda\coloneqq\{k\mid y\in B_{r_{y_{k}}}(y_{k})\}$ and observe that $f_{i-1}(y)\in\{f_{d}(y)\}\cup\{f_{d}(y_{k})\mid k\in\Lambda\}$ and $y\in B_{r_{y_{k}}}(y_{k})\cap B_{r_{y_{i}}}(y_{i})\neq\varnothing$.
    In addition, $y\in B_{r_{y_{i}}}(y_{i})$, and therefore $f_{d}(y)\in cl(f_{d}(y_{i}))$ by the hypothesis on $r_{y_{i}}$. Thus, given that the closure structure on $X\subset R(X)$ is symmetric, it follows that $f_{d}(y_{i})\in cl(f_{i-1}(y))$. Putting this together and observing that $f_{i-1}(y)=f_{i}(y)$ by definition, we have
    $f_{i}(B_{r}(y))\subset f_{i-1}(B_{r}(y))\cup\{f_{d}(y_{i})\}\subset cl(f_{i-1}(y))\cup\{f_{d}(y_{i})\}\subset cl(f_{i-1}(y))=cl(f_{i}(y)).$
  3. Case 3:
    $y\notin\overline{B_{r_{y_{i}}/2}(y_{i})}$. Then there exists $r>0$ such that $B_{r}(y)\subset Y\backslash\overline{B_{r_{y_{i}}/2}(y_{i})}$. Thus $f_{i}(B_{r}(y))=f_{i-1}(B_{r}(y))\subset cl(f_{i-1}(y))=cl(f_{i}(y))$.
Finally, we prove the existence of a homotopy between $f_{i}$ and $f_{i-1}$. Define $H_{i}:Y\times I\rightarrow X\subset R(X)$ such that
$H_{i}(y,t)\coloneqq\left\{\begin{array}[]{ll}f_{i}(y)&t=1\\ f_{i-1}(y)&t\in[0,1)\end{array}\right.$
for $i\in\{1,\ldots,n\}$. Observe that for $t\in[0,1)$ we can find $r>0$ small enough such $H_{i}(B_{r}(y,t))=f_{i-1}(B_{r}(y,t))$, and therefore $H$ is continuous on $Y\times[0,1)$.
We now show that $H$ is continuous at any point of the form $(y,1)\in Y\times[0,1]$, considering the same three cases as above:
  1. Case 1:
    $y\in B_{r_{y_{i}}/2}(y_{i})$. Define $\Lambda\coloneqq\{k\mid y\in B_{r_{y_{k}}/2}(y_{k})\}$, then
    $f_{d}(y)\in f(B_{r_{y_{k}}/2}(y_{k}))\subset cl(f_{d}(y_{k}))$
    for every $k\in\Lambda$ and there exists $r>0$ such that $f_{d}(B_{r}(y))\subset cl(f_{d}(y))$ and $B_{r}(y)\subset B_{r_{y_{k}}/2}(y_{k})$ for every $k\in\Lambda$. Recall that, $\forall y\in Y$, $r_{y}>0$ was chosen so that $f_{d}(B_{r_{y}}(y))\subset cl(f_{d}(y))$. Therefore, $B_{r}(y)\subset B_{r_{y_{i}}/2}$ implies that $f_{d}(B_{r}(y))\subset cl(f_{d}(y_{i}))$. Also, by construction, $B_{r_{y_{k}}}(y_{k})\cap B_{r_{y_{i}}}(y_{i})\neq\emptyset$ for all $k\in\Lambda$, and from the observations above, it follows that $f_{d}(y_{k})\in cl(f_{d}(y_{i}))$ for all $k\in\Lambda$. This implies that
    $f_{i-1}(B_{r}(y))\subset$ $f_{d}(B_{r}(y))\cup\{f_{d}(y_{j})\mid j\in\Lambda\}\subset$
    $cl(f_{d}(y_{i}))\cup\{f_{d}(y_{j})\mid j\in\Lambda\}\subset cl(f_{d}(y_{i})).$
    Furthermore, $f_{i}(y)=f_{d}(y_{i})$ and $f_{i}(B_{r}(y))=\{f_{d}(y_{i})\}$ by the definition of $f_{i}$ and the choice of $r>0$. We conclude that
    $H_{i}(B_{r}(y,1))\subset f_{i}(B_{r}(y))\cup f_{i-1}(B_{r}(y))\subset cl(f_{d}(y_{i}))=cl(f_{i}(y))=cl(H_{i}(y,1)),$
    as desired.
  2. Case 2:
    $y\in\partial B_{r_{y_{i}}/2}(y_{i})$. Then there exists $r>0$ such that $f_{i}(B_{r}(y))\subset cl(f_{i}(y))$ and $f_{i-1}(B_{r}(y))\subset cl(f_{i-1}(y))=cl(f_{i}(y))$. Concluding $H_{i}(B_{r}(y,1))\subset f_{i}(B_{r}(y))\cup f_{i-1}(B_{r}(y))\subset cl(f_{i}(y))$.
  3. Case 3:
    $y\notin\overline{B_{r_{y_{i}}/2}(y_{i})}$. Then there exists $r>0$ such that
    $H_{i}(B_{r}(y,1))=$ $f_{i}(B_{r}(y))=f_{i-1}(B_{r}(y))\subset$
    $cl(f_{i-1}(B_{r}(y))=cl(f_{i}(B_{r}(y))=cl(H_{i}(B_{r}(y,1))).$
Since we replace the values in the image of $B_{r_{y_{i}}/2}(y_{i})$ by $y_{i}$, then $f_{n}$ has just a finite number of value in its image. It now follows that $f^{\prime}_{d}\coloneqq f_{n}$ satisfies the conclusion of the lemma. ∎
In 3.6 and 3.2 below, we have that the domain is comapact and connected so we can apply 2.14, which is all we will use in the remainder of the article; however, we keep 2.13 above to provide a different sufficient conditionfor the image of the discrete modification to be finite which may be useful in the future.
The maps $f_{i}$ in the proof of 2.14 are an example of the following definition: The $v$-flood; this map is an adaption from the $v$-flood in digital topology [Lupton2023DigitalHom].
Definition 2.15 (The $v$-flood).
Given a continuous map $f:Y\rightarrow R(X)$ and $v\in X\subset R(X)$, and let $B\subset Y$. We define the not-necessarily continuous function $f^{v}_{B}:Y\rightarrow R(X)$
$f^{v}_{B}(y)\coloneqq\left\{\begin{array}[]{ll}v&y\in B\\ f(y)&\text{otherwise}\end{array}\right.$
which we call the flood of $f$ on $B$ with $v$ or the $v$-flood of $f$ on $B$. (See Figure 2.)
Figure 2. Illustration of a flood. The pictures represent the image in a region of $Y$ with blue, green and red representing point in $X$; in particular $v$ is red. The first picture represents the iamge of $f$. In the second picture the gray transparent zone is $B$ in that region. The last picture represents the iamge of $f_{B}^{v}$.
Lemma 2.16.
Let $Y$ be a metric space with the topology induced by the metric. Let $f:Y\rightarrow R(X)$ be a function such that $Im(f)\subset X\subset R(X)$ and let $v\in X\subset R(X)$. If for every $y\in Y$ such that $f(y)=v$, there exists an $r_{y}>0$ such that $f(B_{r_{y}}(y))\subset cl(v)$, then $f_{B}^{v}$ is continuous in $X\subset R(X)$ and $f\simeq f_{B}^{v}$ for
$B\coloneqq\bigcup_{y\in f^{-1}(v)}B_{r^{\prime}_{y}/2}(y)$
and any $r^{\prime}_{y}\leq r_{y}$.
Proof.
First we show continuity of $f^{v}_{B}$. We consider two cases:
  1. (1)
    Let $y\in B\subset Y$. Since $B$ is open in $Y$, being the union of open sets, there exists $r>0$, in particular $r_{y}$, such that $B_{r}(y)\subset B$ and therefore $f_{B}^{v}(B_{r}(y))=v\in cl(v)$. In particular, $f_{B}^{v}(B_{r}(y))\subset cl(v)$.
  2. (2)
    Now let $y\in Y\backslash B$. Then by continuity of $f$, and being $cl(f(y))$ the minimal neighborhood of $f(y)$ in $R(X)$, there exists $r>0$ such that $f(B_{r}(y))\subset cl(f(y))$. Suppose that:
    1. (a)
      $v\in cl(f(y))$. Then $f_{B}^{v}(B_{r}(y))\subset f(B_{r}(y))\cup\{v\}\subset cl(f(y))$.
    2. (b)
      $v\notin cl(f(y))$. Since $cl(f(y))$ is the minimal neighborhood, there exists $r^{\prime}>0$ such that $f(B_{r^{\prime}}(y))$ does not contain which is not neighbor of $f(y)$, in particular $v\notin f(B_{r^{\prime}}(y))$. That $r^{\prime}$ has to be less or equal to the distance from $y$ to any point with image $v$ under $f$; thus $B_{\min\{r,r^{\prime}\}/2}(y)\subset Y\backslash B$ and $f^{v}(B_{\min(r,r^{\prime})/2}(y))=f(B_{\min(r,r^{\prime})/2}(y))$. Therefore $f_{B}^{v}(B_{\min(r,r^{\prime})/2}(y))=f(B_{\min(r,r^{\prime})/2}(y))\subset cl(f(y))=cl(f^{v}(y))$
The next step is to define the homotopy. We define $H:Y\times I\rightarrow X\subset R(X)$ by
$H(y,t)\coloneqq\left\{\begin{array}[]{ll}f(y)&\text{ if }t\in[0,1)\\ f_{B}^{v}(y)&\text{ if }t=1\end{array}\right.$
Analogous to 2.12, observe that $H$ is continuous in $Y\times[0,1)$. Now, consider a point $(y,1)$. We have have two cases:
  1. (1)
    If $f(y)=f^{v}(y)$, then there exist $r,r^{\prime}>0$ such that $f(B_{r}(y))\subset cl(f(y))=cl(f^{v}(y))$ and $f_{B}^{v}(B_{r^{\prime}}(y))\subset cl(f^{v}(y))$, thus $H(B_{\min(r,r^{\prime})}(y,1))\subset f(B_{r}(y))\cup f_{B}^{v}(B_{r^{\prime}}(y))\subset cl(f^{v}(y))=cl(H(y,1))$.
  2. (2)
    If $f(y)\neq f^{v}(y)$, then $f_{B}^{v}(y)=v$ and there exists $r>0$ and $y_{0}\in f^{-1}(v)$ such that $f(B_{r}(y))\subset f(B_{r_{y_{0}}}(y_{0}))\subset cl(v)$. For the same reason, there exists $r^{\prime}>0$ such that $f_{B}^{v}(B_{r^{\prime}}(y))=\{v\}$, then $f_{B}^{v}(B_{\min(r,r^{\prime})}(y))=f(B_{\min(r,r^{\prime})}(y))\cup\{v\}\subset cl(v)$, and $H$ continuous.∎
Later when we work with homotopy groups, then we need a version of 2.16 which conserve a point. We prove the extension of this result for preserving a compact set.
Lemma 2.17.
Let $Y$ be a metric space with topology induced by the metric and $A\subset Y$ compact, and let $v,v^{\prime}\in X\subset R(X)$. Let $f:(Y,A)\rightarrow(R(X),v^{\prime})$ be a function such that $Im(f)\subset X\subset R(X)$. If for every $y\in Y$ such that $f(y)=v$, there exists $r_{y}>0$ such that
  • $f(B_{r_{y}}(y))\subset cl(v)$ and
  • if $v\neq v^{\prime}$, $A\cap B_{r_{y}}(y)=\varnothing$,
then $f_{B}^{v}(A)=v^{\prime}$, $f_{B}^{v}$ is continuous in $X\subset R(X)$ and $f\simeq f_{B}^{v}$ rel($A$) in the closure space $R(X)$ for
$B\coloneqq\bigcup_{y\in f^{-1}(v)}B_{r^{\prime}_{y}/2}(y)$
with $r^{\prime}_{y}\leq r_{y}$.
Proof.
First we prove that $f_{B}^{v}(A)=v^{\prime}$. If $v=v^{\prime}$ there is nothing to do, by hypothesis $f_{B}^{v}(f^{-1}(v))=v$. On the other hand, if $v\neq v^{\prime}$, then $B\subset Y\backslash A$. Thus $f^{v}_{B}(A)=f(A)=v^{\prime}$ by definition.
Observe that $f_{B}^{v}$ is continuous and $f\simeq f_{B}^{v}$ for 2.16. Also in 2.16, we define the homotopy such that $H(-,1)=f_{B}^{v}$ and $H(-,t)=f$ for every $t\in[0,1)$. Thus $H(A,t)=v^{\prime}$, and then the homotopy is relative to $A$. ∎
Given a compact metric space $Y$ and $A\subset Y$ compact, for any map $f:Y\rightarrow R(X)$ such that $f(A)=v^{\prime}\in X$, we have that in the image of $\Theta_{X}f:Y\rightarrow X\subset R(X)$ there is just a finite bunch and $\Theta_{X}f(A)=v^{\prime}$ by 2.13. In addition, for every $y\in Y$ such that $f(y)=v_{0}$, by continuity and $A$ being a compact, we can find $r_{y}>0$ such that $f(B_{r_{y}}(y))\subset cl(v_{0})$ and $A\cap B_{r_{y}}(y)=\varnothing$. Thus, by 2.17, we conclude that $(\Theta_{X}f)^{v_{0}}_{B_{0}}\simeq\Theta_{X}f\simeq f$ and $(\Theta_{X}f)^{v_{0}}_{B_{0}}(A)=v^{\prime}$ with
$B_{0}=\bigcup_{y\in(\Theta_{X}f)^{-1}(v_{0})}B_{r_{y}/2}(y)$
Observe we can repeat this process for $(\Theta_{X}f)^{v_{0}}_{B_{0}}$, $v_{1}$ and building in a similar fashion $B_{1}\subset Y$; obtaining $((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}}$ which is homotopy to $(\Theta_{X}f)^{v_{0}}_{B_{0}}$ and preserves the image in $A$.
Then we repeat this process a finite amount of times for all of the vertices in $Im(\Theta_{X}f)$, obtaining
$f\simeq$ $(\Theta_{X}f)^{v_{0}}_{B_{0}}\simeq((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}}\simeq\ldots\simeq(\cdots((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}})\cdots)_{B_{n}}^{v_{n}}$
$v^{\prime}=$ $(\cdots((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}})\cdots)_{B_{n}}^{v_{n}}(A)$
. For simplification, we denote $(\cdots((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}})\cdots)_{B_{n}}^{v_{n}}$ by $f^{v_{n}}$.
Proposition 2.18.
Suppose that we have the sequences $\{y_{n}\}\rightarrow x$ and $\{z_{n}\}\rightarrow x$ in a metric space $X$ and we have a function $f:X\rightarrow Y$ (not necessarily continuous) such that #$\#Y<\infty$. Then we can take a subsequence of $\{(y_{n},z_{n})\}$ such that $f(y_{n})=f(y_{k})$ and $f(z_{n})=f(z_{k})$ for every $k,n$.
Proof.
[…]
Lemma 2.19.
Let $Y$ be a compact metric space and $X$ a graph. Given a continuous map $f:Y\rightarrow R(X)$ with image $\{v_{0},\ldots,v_{m}\}$, for every $x\in Y$ there exists $r_{x}>0$ such that if $v,w\in f^{v_{m}}(B_{r_{x}}(x))$, then $v\in cl(w)$ (equivalently $w\in cl(v)$).
Proof.
We proceed by contradiction. Suppose that there exists $x\in Y$ such that for every $r>0$ there exists a pair $w_{r},w^{\prime}_{r}\in f^{v_{m}}(B_{r}(x))$ we have that $w_{r}\notin cl(w^{\prime}_{r})$. We can restrict just to the elements in $1/k$ for every $k\in\mathbb{N}$; in consequence we have elements $y_{k},z_{k}\in X$ such that $y_{k},z_{k}\in B_{1/k}(x)$, $f^{v_{m}}(y_{k})=w_{1/k}$ and $f^{v_{m}}(z_{k})=w^{\prime}_{1/k}$. Thus we have the sequences $\{y_{k}\}$ and $\{z_{k}\}$ converge to $x$; therefore, by 2.18, we can take a subsequence of both such that the image of all of the elements of the subsequence $\{y_{k}\}$ (resp. $\{z_{k}\}$) is a fixed element $w$ (resp. $w^{\prime}$).
We name $v\coloneqq f^{v_{m}}(x)$, and divide this proof in two cases:
  1. (1)
    Suppose that $w<v$ or $w^{\prime}<v$, we obtain a direct contradiction. The elements only change their values to $w$ (resp. $w^{\prime}$) only appears at $f^{w}$ (resp. $f^{w^{\prime}}$), and in $f^{v}(B_{r}(x))=v$ for some $r>0$. (See Figure 3.)
    Figure 3. Illustration of $w<v$ or $w^{\prime}<v$. The pictures represent the image in a region of $Y$ with blue, green and red representing $w$, $w^{\prime}$ and $v$ respectively. The white point with black contour is $x$. From left to right: The first image is before $f^{v}$. Second image represents a point with image $v$ and its neighborhood. The last one represents $f^{v}$. (We just ignored all of the rest of points in the image.)
  2. (2)
    Suppose that $w>v$ and $w^{\prime}>v$. Without loss of generality, we can assume that $w<w^{\prime}$. Observe once again that in the map $f^{v}$ we have a $r>0$ such that the image of $B_{r}(x)$ is $v$. Then in $f^{v}$ the image of $\{y_{k}\}$ must become $w$. Take any element $z\in S^{m}$ such that $f^{w}(z)=w^{\prime}$, then
    $\lim_{n}d(y_{k},z)\rightarrow d(x,z),$
[…]
    Figure 4. Illustration of $v<w<w^{\prime}$. The pictures represent the image in a region of $Y$ with blue, green and red representing $w$, $w^{\prime}$ and $v$ respectively. The white point with black contour is $x$. From left to right: The first image represent $f^{w}$, and then we must have a sequence with image $w$; we have a green point in the contour of the red circle. In the second image, the circles the distance and the half of the distance between the green and the white point $x$. The last one represents $f^{w^{\prime}}$ and shows the space without green points. (We just ignored all of the rest of points in the image.)
Observe that $v=w$ or $v=w^{\prime}$ is also an automatic contradiction. ∎
The discrete modification (2.10) is built to convert all of the point in the image into vertexes. The next definition seeks to perform an inverse process, moving from a function with only image at the vertices to one with image at $|VR(X)|$.
Definition 2.20 (The Convex Transformation).
Let $X$ be a graph, be $Y$ a triangulizable compact space (with a finite amount of triangles), $\mathcal{U}$ an open cover of $Y$ and $f:Y\rightarrow X\subset R(X)$ such that:
  1. (1)
    For every $U\in\mathcal{U}$, all $y,y^{\prime}\in U$ satisfies that $f(y)\in c_{X}(f(y^{\prime}))$.
  2. (2)
    Every triangle in $Y$ is totally contained in some $U\in\mathcal{U}$.
We define the convex transformation of $f$, denoted by $f^{\textbf{conv}}:Y\rightarrow|VR(X)|$, such that for every triangle in the triangulation of $Y$ we look at its vertices, say $\{y_{0},\ldots,y_{n}\}$, and define
$f^{conv}\left(\sum_{i=0}^{n}t_{i}y_{i}\right)=$ $\sum_{i=0}^{n}t_{i}f(y_{i})$
such that $t_{i}\geq 0\text{ and }\sum_{i=0}^{n}t_{i}=1$
Lemma 2.21.
Consider $X$, $Y$, $\mathcal{U}$ and $f$ as in 2.20. $f^{\textbf{conv}}:Y\rightarrow(|VR(X)|,c_{|VR(X)|})$ is well-defined
Proof.
To show that $f^{\textbf{conv}}$ is well-defined, it is enough to show that if of two simplexes, say $\Delta^{n}=\{v_{0},\ldots,v_{n}\}$ and $\Delta^{m}=\{w_{0},\ldots,w_{m}\}$, share a face, say $\Delta^{n}=\{x_{0},\ldots,x_{n}\}$ such that $x_{i}=v_{i}=w_{i}$ for $i\in[k]$, then the value $f^{\textbf{conv}}$ is exactly the same in shared face. For $t_{i},s_{j},r_{a}\in[0,1]$ for every $i\in[n],j\in[m],a\in[k]$ such that $\sum_{i=0}^{n}t_{i}=\sum_{j=0}^{m}s_{j}=\sum_{a=0}^{k}k_{a}=1$ and $t_{i}=s_{j}=0$ for $i\in[n]\backslash[k],j\in[n]\backslash[k]$ we that
$f^{\textbf{conv}}\mid_{\Delta^{k}\subset\Delta^{n}}\left(\sum_{i=0}^{n}t_{i}x_{i}\right)=$ $\sum_{i=0}^{n}t_{i}f^{v_{m}}(v_{i})=\sum_{a=0}^{k}r_{a}f^{v_{m}}(x_{a})=f^{\textbf{conv}}\mid_{\Delta^{k}}\left(\sum_{a=0}^{k}r_{a}x_{a}\right)$
$=$ $\sum_{j=0}^{m}s_{j}f^{v_{m}}(w_{j})=f^{\textbf{conv}}\mid_{\Delta^{k}\subset\Delta^{m}}\left(\sum_{j=0}^{m}s_{j}w_{j}\right).$
Obtaining that $f^{\textbf{conv}}$ is well-defined. ∎
Lemma 2.22.
Consider $X$, $Y$, $\mathcal{U}$ and $f$ as in 2.20. $f^{\textbf{conv}}:Y\rightarrow(|VR(X)|,c_{|VR(X)|})$ is continuous and $f^{\textbf{conv}}\simeq f$ in $R(X)$.
Proof.
By construction, it follows that $f^{conv}$ is continuous in the triangles because that map is just the geometric realization from an $n$-simplex. Now, since we have a locally finite closed cover, observe that $f^{conv}$ is continuous by 17 A.18 in [Cech_1966].
The next step is to build the homotopy. Let $H:S^{n}\times I\rightarrow R(X)$ be the map
$H(x,t)\coloneqq\left\{\begin{array}[]{ll}f^{conv}(x)&\text{ if }t\in[0,1)\\ f(x)&\text{ if }t=1.\end{array}\right.$
As we have seen before, continuity in $S^{n}\times[0,1)$ is straightforward: we can take a radius small enough such that $B_{r}(x,t)\cap S^{n}\times\{I\}=\varnothing$, and then $H(B_{r}(x,t))=B_{r}(x,t)$; thus, by continuity in $f^{\textbf{conv}}$, we obtain the continuity in that set. Consider a point $(x,1)$; by construction, the image of every triangle $\Delta$ in $Y$ is inside an open set $U$ such that for all $y,y^{\prime}\in U$ satisfies that $f(y)\in cl(f(y^{\prime}))$. Thus, every point $z\in f(\delta)$ is neighbor of the vertexes as well as the elements of the geometric realization of that triangle (a.k.a $f^{conv}$ of them). Therefore, remembering that $f(x)\in X\subset R(X)$, for every triangle $\{\Delta_{1},\ldots,\Delta_{a}\}$ such that $x\in\Delta_{i}$ we have that $\cup_{i=1}^{a}f^{\textbf{conv}}(\Delta_{i})\subset cl(f(x))$, and then for $r>0$ such that $B_{r}(f(x))\subset\cup_{i=1}^{a}\Delta_{i}$, we obtain that $H(B_{r}(x,1))\subset f^{\textbf{conv}}(x)\cup f(x)\subset cl(f(x))$. ∎
3. Weak Homotopy Equivalence of $(X,\lambda)$ and $|VR(X,\lambda)|$
In this section, we prove that the homomorphism $\pi_{n}(1_{|VR(X)|}):\pi_{n}(|VR(X)|)\rightarrow\pi_{n}(R(X))$ is an isomorphism, with $1_{|VR(X)|}:|VR(X)|\rightarrow R(X)$ the inclusion map. The core of the proof is seeing some kind of functions to $R(X)$ are homotopic to maps which image in $X\subset R(X)$ satisfying that for every point in the domain there exist a neighborhood $U$ such that every pair of point in $U$ are neighbors; and then we can prove that the later map is homotpic to some realization of a simplicial map.
We define $S^{n}\coloneqq\{x\in\mathbb{R}^{n+1}\mid|x|=1\}$ with its topological structure and metric inherited from $\mathbb{R}^{n+1}$.
We need a last classical technical result before starting the proof of $\pi_{n}(1_{|VR(X)|})$ is an epimorphism. We leave this result without proof.
Lemma 3.1 (Lebesgue’s number lemma).
Let $X$ be a compact metric space and $\mathcal{U}$ an open cover of $X$, then there exists a number $\delta>0$ such that every subset of $X$ having diameter less than $\delta$ is contained in some member of the cover.
Proposition 3.2.
Let $X$ be a graph. Every map $f:S^{n}\rightarrow R(X)$ is homotopic to the composition of $1_{|VR(X)|}:|VR(X)|\rightarrow R(X)$ and some map $g:S^{n}\rightarrow|VR(X)|$. This implies that $\pi_{n}(1_{|VR(X)|})$ is an epimorphism.
Proof.
Let $f:(S^{n},*)\rightarrow(R(X),v)$ a continuous map with $v\in X$. Since $S^{n}$ is compact and connected, $\Theta_{X}f\simeq f$ in $R(X)$ and $Im(\Theta_{X}f)$ is a finite amount of vertexes, say $\{v_{0},\ldots,v_{m}\}$, by 2.14.
[…]
With similar definitions, we can build $B_{1},\ldots,B_{m}$ and obtain $f^{v_{m}}$, observing that
$f\simeq$ $(\Theta_{X}f)^{v_{0}}_{B_{0}}\simeq((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}}\simeq\ldots\simeq(\cdots((\Theta_{X}f)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}})\cdots)_{B_{m}}^{v_{m}}\eqqcolon f^{v_{m}}$
$v=$ $f^{v_{m}}(*)$
Thus, by 2.19, for every $x\in S^{n}$, there exists $r>0$ such that for every $w,w^{\prime}\in f^{v_{m}}(B_{r}(x))$ we have that $w\in cl(w^{\prime})$; we denote this open cover of $S^{n}$ as $\mathcal{U}$.
Take the classical triangulation of $S^{n}$ such that * is one of the vertexes of its triangles and apply the barycentric subdivision until we have that every triangle has diameter less than the Lebesgue’s number of the cover $\mathcal{U}$. Then, by 3.1 and 2.19, it satisfies the hypothesis of 2.22; thus, $(f^{v_{m}})^{\textbf{conv}}\simeq f^{v_{m}}$, and $(f^{v_{m}})^{\textbf{conv}}$ is continuous with the closure operation $c_{|VR(X)|}$.
In conclusion, $f\simeq f^{v_{m}}\simeq(f^{v_{m}})^{\textbf{conv}}$, and the latter is continuous in the geometric realization of the Vietoris-Rips complex. ∎
To prove $\pi_{n}(1_{|VR(X)|})$ is a monomorphism, we are going to need two more technical results. Following the notation of [Spanier_1966], we denote as sd the barycentric subdivision and $\text{sd}^{n}$ when we apply the map $n$-times.
Definition 3.3.
Let $K_{1}$ and $K_{2}$ be simplicial complexes and let $f:|K_{1}|\rightarrow|K_{2}|$ be a continuous map. A simplicial map $\varphi:K_{1}\rightarrow K_{2}$ is called a simplicial approximation to $f$ if $f(\alpha)\in|s_{2}|$ implies that $|\varphi|(\alpha)\in|s_{2}|$ for $\alpha\in|K_{1}|$ and $s_{2}\in K_{2}$.
Theorem 3.4 (Simplicial-approximation Theorem; [Spanier_1966], 3.5.8).
Let $(K_{1},L_{1})$ be a finite simplicial pair and let $f:(|K_{1}|,|L_{1}|)\rightarrow(|K_{2}|,|L_{2}|)$ be a map. There exists an integer $N$ such that if $n\geq N$ there are a simplicial approximation $(\text{sd}^{n}K_{1},\text{sd}^{n}L_{1})\rightarrow(K_{2},L_{2})$ to $f$.
Thus, considering for example the canonical triangulation of $S^{n}$, the realization of simplicial maps characterizes all of the maps from $(S^{n},*)$ to $(|VR(X)|,v)$.
The following results claims that the convex modification is homotopic (with the closure operation $c_{|VR(X)|}$) to the convex modification of the barycentric subdivision. This result is essential in the proof of having a monomorphism and implies some technical conditions in the selection of balls when we apply the floods in that proof.
Proposition 3.5.
Let $Y$ a triangulizable compact metric space (with finite triangulation, say $K_{1}$) and $f:Y\rightarrow X$ such that every triangle in the triangulation is within a simplex. Then $f^{\textbf{conv}}$ and $(\text{sd}f)^{\textbf{conv}}:|\text{sd}K_{1}|\rightarrow|VR(X)|$ are homotopy equivalent in $|VR(X)|$.
Proof.
Consider an $n$-simplex $\Delta=\{x_{0},\ldots,x_{n}\}$ in the triangulation such that $x_{0}<x_{1}<\ldots<x_{n}$ and describe the interval $I$ as the geometric realization of $\{z_{0},z_{1}\}$. Recall that $|\Delta|\times I$ is a triangulizable metric space with $(n+1)$-simplexes such that after any $(w_{i},w^{\prime}_{i})\in\Delta\times\{z_{0},z_{1}\}$ such that either $w_{i}<w_{i+1}$ and $w^{\prime}_{i}=w^{\prime}_{i+1}$ or $w_{i}=w_{i+1}$ and $w^{\prime}_{i}<w^{\prime}_{i+1}$. We work with $\Delta\in\text{sd}K_{1}$.
Define $f^{conv}$ as the convex modification of $f$. Let $\Delta$ be an $n$-simplice in $\text{sd}K_{1}$, say
$\left\{x_{0},\frac{1}{2}(x_{0}+x_{1}),\frac{1}{3}(x_{0}+x_{1}+x_{2}),\ldots,\frac{1}{n+1}\left(\sum_{i=0}^{n}x_{i}\right)\right\}$
(interchanging the $x_{i}$’s, you actually have all of the simplices in $\text{sd}K_{1}$). Then take $f^{\textbf{conv}}\mid_{|\Delta|}$ and define $(\text{sd}f)^{\textbf{conv}}\mid_{|\Delta|}:|\Delta|\rightarrow|VR(X)|$ such that
$(\text{sd}f)^{\textbf{conv}}\mid_{|\Delta|}\left(\sum_{i=0}^{n}t_{i}x_{i}\right)\coloneqq\sum_{i=0}^{n}t_{i}f(x_{i})$
Now, we can triangulate $|\Delta|\times I$ and define $H:|\Delta|\times I\rightarrow|VR(X)|$ such that for every $n+1$-simplex $\{(w_{0},w^{\prime}_{0}),\dots,(w_{n+1},w^{\prime}_{n+1})\}$
$H\left(\sum_{i=0}^{n+1}t_{i}(w_{i},w^{\prime}_{i})\right)\coloneqq\sum_{i=0}^{n+1}t_{i}H(w_{i},w^{\prime}_{i})$
with $H(w_{i},w^{\prime}_{i})=f^{\textbf{conv}}(w_{i})$ if $w^{\prime}_{1}=z_{0}$ and $H(w_{i},w^{\prime}_{i})=(\text{sd}f)^{\textbf{conv}}(w_{i})\mid_{|\Delta|}$ if $w^{\prime}_{1}=z_{1}$.
We might join every simplex to extend these maps to $|K_{1}|\rightarrow|VR(X)|$ and by construction we obtain that $H(y,0)=f^{\textbf{conv}}(y)$ and $H(y,1)=(\text{sd}f)^{\textbf{conv}}(y)$; concluding that me have the homotopy between the both maps. ∎
Proposition 3.6.
Let $X$ be a graph. $\pi_{n}(1_{|VR(X)|}):\pi_{n}(|VR(X)|)\rightarrow\pi_{n}(R(X))$ is a monomorphism.
Proof.
Let $f:(S^{n},*)\rightarrow(|VR(X)|,v)$ such that $1_{|VR(X)|}f:(S^{n},*)\rightarrow(R(X),v)$ is homotopic to the constant map $\hat{v}:(S^{n},*)\rightarrow(R(X),v)$ such that $\hat{v}(x)=v$ for every $x\in S^{n}$, i.e. there exists $H:S^{n}\times I\times R(X)$ such that $H(\{*\}\times I)=v$, $H(x,0)=f(x)$ and $H(x,1)=\hat{v}(x)=v$. We want to transform $H$ to some continuous map in $|VR(X)|$ with the same properties.
Without loss of generality, we can consider $f$ as a simplicial-approximation to itself by 3.4; this is because we only look for a homotopy, and then we do not have problems if we start with a map homotopic (in $|VR(X)|$) to the original.
In the same sense of 3.2, because $S^{n}\times I$ is compact, then $\Theta_{X}H:S^{n}\times I\rightarrow X\subset R(X)$ satisfies that $\Theta_{X}H\simeq H$ and #$\#Im(\Theta_{X}H)<\infty$, say $\{v_{0},\ldots,v_{m}\}$.
Only in this definition, we say that:
  • $(x,t)$ is in the bottom (of $S^{n}\times I$) if $t=0$.
  • $(x,t)$ is in the top (of $S^{n}\times I$) if $t=1$.
  • $(x,t)$ is in the middle (of $S^{n}\times I$) if $t\in(0,1)$.
Since $\Theta_{X}H$ is continuous in $R(X)$ and {}×$\{*\}\times I$ is compact, we have that for every $(x,t)\in S^{n}\times I$ such that $H(x,t)=v_{0}$ there exists $r_{(x,t)}>0$ such that $\Theta_{X}H(B_{r_{(x,t)}}(x,t))\subset cl(H(x,t))$, if $v\neq v_{0}$ then $B_{r_{(x,t)}}(x,t)\cap(\{*\}\times I)=\varnothing$, and
  • if is in the top, $B_{r_{(x,t)}}(x,t)\cap S^{n}\times\{0\}=\varnothing$.
  • if is in the middle, $B_{r_{(x,t)}}(x,t)\cap S^{n}\times\{0,1\}=\varnothing$.
  • if is in the bottom, $B_{r_{(x,t)}}(x,t)\cap S^{n}\times\{1\}=\varnothing$ and $x$ does not touch any vertex or face of the triangulation of $S^{n}$ but itself, this is possible because we have just a finite amount of triangles.
Thus, by 2.17, we obtain that $(\Theta_{X}H)_{B_{0}}^{v_{0}}\simeq\Theta_{X}H$, $(\Theta_{X}H)_{B_{0}}^{v_{0}}(\{*\}\times I)=v$, $(\Theta_{X}H)_{B_{0}}^{v_{0}}(x,1)=v$ and it does not change the value of the vertexes on the triangulation of $S^{n}$; defining $B_{0}$ as the union of those balls with center $(x,t)\in(\Theta_{X}H)^{-1}(v)$.
With similar definitions, we can build $B_{1},\ldots,B_{m}$ and obtain $H^{v_{m}}$, observing that
$H\simeq$ $(\Theta_{X}H)^{v_{0}}_{B_{0}}\simeq((\Theta_{X}H)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}}\simeq\ldots\simeq(\cdots((\Theta_{X}H)^{v_{0}}_{B_{0}})^{v_{1}}_{B_{1}})\cdots)_{B_{m}}^{v_{m}}\eqqcolon H^{v_{m}}$
$v=$ $H^{v_{m}}(*)$
Thus, by 2.19, for every $(x,t)\in S^{n}\times I$ there exists $r>0$ such that for every $w,w^{\prime}\in H^{v_{m}}(B_{r}(x,t))$ we have that $w\in cl(w^{\prime})$; we denote this open cover of $S^{n}\times I$ as $\mathcal{U}$.
Take the triangulation of the simplicial approximation $K_{1}$ and describe $I$ as the geometric realization of $\{0,1\}$. Now we take the barycentric subdivision in both of them until the diameter of the simplices in $\text{sd}^{k}K_{1}\times\text{sd}^{k}\{0,1\}$ is less than the Lebesgue’s number of $\mathcal{U}$. Then, by 3.1 and 2.19, these objects satisfies the hypothesis of 2.22. Thus $(H^{v_{m}})^{\textbf{conv}}$ is continuous in $|VR(X)|$, $(H^{v_{m}})^{\textbf{conv}}(x,1)=\hat{v}(x)=v$ and $(H^{v_{m}})^{\textbf{conv}}(x,0)=(\text{sd}^{k}f^{v_{m}})^{\textbf{conv}}(x)$.
Lastly, since the points of the vertexes of the triangulation was not touched, we have that
$(\text{sd}^{k}f^{v_{m}})^{\textbf{conv}}\simeq(\text{sd}^{k-1}f^{v_{m}})^{\textbf{conv}}\simeq\ldots\simeq(\text{sd}f^{v_{m}})^{\textbf{conv}}\simeq(f^{v_{m}})^{\textbf{conv}}=f$
in $|VR(X)|$ by 3.5. Concluding that $f\simeq\hat{v}$ in $|VR(X)|$. ∎
Theorem 3.7.
Let $G=(X,E)$ be a graph, and let $(X,c_{G})$ be the induced Čech closure space. Then $|VR(X)|$ has the same weak homotopy type as $(X,c_{G})$.
Proof.
Given the strong homotopy equivalence in 2.12 with the pair of maps $1_{X}:X\rightarrow R(X)$ and $\Theta_{X}:R(X)\rightarrow X$, we obtain that $\pi_{n}(\Theta_{X}):\pi_{n}(R(X))\rightarrow\pi_{n}(X)$ is an isomorphism.
In addition, for 3.2 and 3.6, we have that $\pi_{n}(1_{|VR(X)|}):\pi_{n}(|VR(X)|)\rightarrow\pi_{n}(R(X))$ is also an isomprhism. Thus, we obtain that $\pi_{n}(\Theta_{X})\pi_{n}(|VR(X)|):\pi_{n}(|VR(X)|)\rightarrow\pi_{n}(X)$ is an isomorphism. ∎
Acknowledgments
I would like to thank my thesis advisor, Antonio Rieser, who has helped with several suggestions about the order of the ideas of these paper to improve the delivery and to reduce the number of particular lemmas for specific cases. I am deeply grateful to the AIM Workshop on Discrete and Combinatorial Homotopy Theory, its organizers: Helen Barcelo, Antonio Rieser and Volkmar Welker, and certainly with my one-week teammates: Greg Lupton, Oleg Musin, Nicholas Scoville and Chris Staecker; I learned a great deal from them about how to decompose a problem, discuss the ideas and try to simplify them in order to improve the exposition. I am grateful to Antonio Rieser for telling me about the existence of [WebJMMNicks] after I showed him the first versions of this work.